# CI/CD Configuration Practices in Open-Source Android Apps: An Empirical Study

### v1

@@ Abstract.
Continuous Integration and Continuous Delivery (CI/CD) is a well-established practice that automatically builds, tests, packages, and deploys software systems. To adopt CI/CD, software developers need to configure their projects using dedicated YML configuration files. Mobile applications have distinct characteristics when it comes to CI/CD practices, such as testing on various emulators and deploying to app stores. However, little is known about the challenges and added value of adopting CI/CD in mobile applications and how developers maintain such a practice. In this paper, we conduct an empirical study on CI/CD practices in $2,564$ Android apps adopting four popular CI/CD services, namely GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD. We observe a lack of commonality and standards across projects and services, leading to complex YML configurations and associated maintenance efforts. We also observe that CI/CD configurations focus primarily on the build setup, with around half of the projects performing standard testing and only 9% incorporating deployment. In addition, we find that CI/CD configurations are changed bi-monthly on average, with frequent maintenance correlating with active issue tracking, project size/age, and community engagement. Our qualitative analysis of commits uncovered 11 themes in CI/CD maintenance activities, with over a third of the changes focusing on improving workflows and fixing build issues, while another third involves updating the build environment, tools, and dependencies. Our study emphasizes the necessity for automation and AI-powered tools to improve CI/CD processes for mobile applications, and advocates for creating adaptable open-source tools to efficiently manage resources, especially in testing and deployment.
@@ 1. Introduction
Continuous Integration and Continuous Delivery (CI/CD) is a well-established practice that automates the building, testing, packaging, and deployment of software systems. (Atlassian, [n.d.]; CircleCI, [n.d.]). CI/CD pipelines consist of automated processes designed to reduce human error, improve developer productivity, and accelerate release cycles. Their widespread adoption in both industry and open-source communities has led to extensive research on the challenges and benefits associated with CI/CD pipelines (Hilton et al., 2017, 2016a; Thatikonda, 2023).
@@
Mobile applications, henceforth referred to as mobile apps, have become integral to our daily lives, serving various purposes, from communication to entertainment. Research reports that apps with more frequent release cycles tend to achieve higher ratings, as users often positively perceive regular updates and improvements (McIlroy et al., 2016). Research has also explored the benefits of adopting CI/CD in the software development lifecycle, notably enhancing quality and accelerating the delivery timeline. In the context of CI/CD, Cruz et al. (2019a) studied the testing practices in open-source mobile apps and found that the use of CI/CD pipelines is surprisingly limited among the mobile apps they analyzed. Mobile applications require distinct CI/CD pipelines compared to other software systems. This is because developers need to ensure extensive testing across multiple emulators and manage deployments to app stores, such as the Google Play Store (Nayebi et al., 2017). Despite such importance, challenges in configuring and maintaining CI/CD for mobile apps have not been well studied. Though Liu et al. (2022) studied the adoption of CI/CD in mobile apps, their study mainly provided general statistical insights about CI/CD usage without exploring in-depth the specific configuration or maintenance practices of developers. In this paper, we conduct an empirical study to investigate the configuration practices and maintenance activities associated with CI/CD in a curated set of 2,564 open-source Android apps adopting four popular CI/CD services, namely GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD. In particular, our study explores the common patterns and challenges that developers face with CI/CD configurations for their apps. We aim to provide insights into the current state of CI/CD adoption within the Android development community and highlight areas for potential improvement. We address the following research questions (RQs).
RQ1: How do CI/CD configurations vary in terms of complexity and tool usage?
@@
In this RQ, we examine CI/CD configurations across multiple services. We observe that CI/CD configurations are not standardized, with complexities varying by service and type of configuration. We also find that most of the attention in CI/CD configuration is on the pipeline setup, accounting for more than 81% of the configurations. While testing is found to be present in about half of the projects, mainly through standard unit tests, deployment is comparatively less common, occurring in 9% of the projects. Our investigation shows that configuring deployment is notably complex and often does not directly target app stores, indicating a need for more integrated and automated solutions to streamline mobile app deployments to popular platforms, such as the Google Play Store.
@@
Adopting CI/CD in mobile app development makes developers responsible for the ongoing maintenance of their pipelines, which involves tasks such as feature additions and issue resolutions. In this RQ, we analyze and model the evolution of CI/CD configurations. We observe that developers regularly maintain CI/CD pipelines, yet not very frequently, mainly focusing on improving the setup and build phases. Over time, projects tend to simplify the CI/CD pipelines for their apps. While maintenance is found to occur bi-monthly on average, many projects update CI/CD pipelines more often, especially those using GitHub Actions. Our results show that increased maintenance activity correlates significantly with factors such as active issue tracking, project size/age, and community engagement.
@@
As CI/CD pipelines are continually updated, it is important to understand the rationale for these changes. In this RQ, we perform a manual analysis of a statistically significant sample of commits related to CI/CD configurations. We observe that more than a third of CI/CD configuration changes focus on improving workflow and fixing build issues, while another third targets updates to the build environment, tools, and dependencies. To validate our manually identified themes for all commits in our dataset, we employ Latent Dirichlet Allocation (LDA) for automated topic generation and observe that the results are consistent with our observations.
@@
  • This study is the first of its kind in analyzing CI/CD configurations in Android apps across various services, using a dataset of 2,564 projects. It highlights current challenges in CI/CD adoption due to non-standardized practices.
  • This study models the evolution of CI/CD configurations using linear regression, and identifies the potentially relevant factors linked to changes in CI/CD pipelines.
  • This study presents a manually validated classification of 11 themes related to the maintenance of CI/CD pipelines, corroborated by automated automatically generated topics using LDA, ensuring both depth and breadth in our analysis.
Overall, our study highlights the lack of standardized configurations in CI/CD pipelines for Android apps, emphasizing the need for uniform practices to improve interoperability. In addition, this study recommends employing automation tools and AI-driven solutions, such as bots and intelligent recommendation systems, to enhance and streamline CI/CD processes. Finally, researchers are encouraged to create adaptable frameworks and open-source tools to address resource constraints in mobile apps, especially those concerned with testing and deployment phases.
Paper organization: The rest of the paper is organized as follows. Section 3 details our study setup to investigate the challenges in configuring CI/CD pipelines for mobile apps. Section 4 outlines our research questions, providing both their motivation and the obtained results. We discuss the implications of our work in Section 5. Section 6 illustrates the potential validity threats of our findings. We present the related work in Section 7. Finally, we conclude our work and provide recommendations for future work in Section 8.
@@ 2. Background
Continuous Integration (CI) and Continuous Delivery (CD) are essential practices in modern software development, particularly for Android mobile applications. These practices ensure that code changes are automatically built, tested, and prepared for release, allowing for faster and more reliable app updates.
Continuous integration allows developers to frequently commit code changes to a shared, central repository, triggering automated build and test processes for smooth integration. This helps catch bugs early, streamline new features, and improve code quality. Continuous delivery extends CI by automating deployment and preparing code changes for release. The deployment process in CI/CD involves moving integrated and tested changes to a production environment. Deployments can be manual or automated, depending on project needs. Automated deployments reduce the risk of human error and ensure consistency across different environments. Many CI/CD service providers (referred to as CI/CD services) are available to streamline the development process for Android applications. Examples include GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD. Each service offers unique features and configurations tailored to various project needs and development workflows. Triggering CI/CD events can be done in several ways. Some services, like GitHub Actions, have native integrations that allow seamless triggering of CI/CD events within the same platform. Other services, like CircleCI, enable CI/CD builds to be triggered on third-party servers using webhooks for specific events, such as when code is pushed or a pull request is created. In this section, we describe the CI/CD process in general, how it is configured in the context of Android mobile app development, and the differences in CI/CD configuration across various services.
@@ 2.1. CI/CD Process
In an Android CI/CD pipeline, various processes are involved to ensure seamless integration and delivery. Figure 1 gives an overview of the CI/CD process in general. These processes typically include Environment Setup, Dependency Management, Compiling/Building, Packaging, Testing, Deployment, and Feedback. In Listing 1, we provide a sample YML configuration for a CI/CD pipeline (in GitHub Actions) and give a brief description of each process. The red text represents the CI/CD directives (keys that define specific instructions) that guide the build, blue text represents executable commands, black text represents parameter values, and green text represents comments.
Figure 1. An overview of our data collection process.
  • Environment Preparation. In this stage, the CI/CD service initializes the environment needed to build and test the Android application, which involves choosing the machine operating system, specifying the appropriate version of the Java Development Kit (JDK), setting up Android SDK components, and configuring any necessary environment variables (see lines 5-12).
  • Dependency Management. In this stage, the dependencies required to build and run the Android application are installed, which typically involves running tasks for the Gradle [^1] https://gradle.org build management tool, for example, to download and cache dependencies specified in build files (see lines 14-15).
  • Building. In this stage, the Android application is compiled by running respective build tasks to compile the source code and generate executables and resources (see lines 17-18).
  • Packaging. In this stage, the Android application is packaged into a certain distribution format, such as the Android Package (APK) or the Android App Bundle (AAB), which comprises executables with any required configurations and third-party libraries (see lines 20-25).
@@
  • Deployment. In this stage, the successfully built and tested application is deployed to a production environment, which typically involves uploading the APK file to a distribution platform, such as the Google Play Store (see lines 30-33).
@@ 2.2.1. GitHub Actions
GitHub Actions is directly integrated within GitHub, allowing developers to define their CI/CD workflows using YML files placed in the .github/workflows directory. This service is highly flexible, enabling developers to customize every step of the CI/CD process using job and step directives. Its integration with GitHub repositories facilitates the triggering of workflows based on various events (e.g., pushes, pull requests, and more). In addition, GitHub Actions provides access to a wide range of pre-built actions available in the GitHub Marketplace, which can be easily incorporated into workflows.
@@ 2.2.2. Travis CI
This is an independent CI/CD service that simplifies the CI/CD setup with its use of a .travis.yml file located in the root of the repository. It employs predefined directives, such as install, script, before_script, and after_success, making it straightforward to set up and use. This allows developers to quickly establish a CI/CD pipeline without needing to define every process explicitly. Travis CI integrates seamlessly with GitHub repositories, making it a popular choice for many open-source projects due to its simplicity and ease of use.
@@ 2.2.3. CircleCI
This is an independent CI/CD service that uses a circle.yml or config.yml file found in the .circleci/ directory. It provides predefined jobs like checkout, run, and deploy, which help structure the CI/CD process while allowing for the inclusion of custom scripts for more complex setups. CircleCI is known for its strong support for Docker, enabling containerized builds that ensure consistency across different environments [^2] https://circleci.com/docs/about-circleci. This makes CircleCI suitable for projects requiring advanced build and deployment configurations, such as Android mobile apps.
@@ 2.2.4. GitLab CI/CD
This is fully integrated within the GitLab version control system, but can also be used within GitHub. It uses a .gitlab-ci.yml file to define the CI/CD pipeline. Similar to Travis CI, it features predefined stages such as before_script, script, and after_script, while also allowing developers to create custom stages and jobs for greater flexibility. The integration with GitHub repositories is straightforward, allowing developers to connect their GitHub projects to GitLab CI/CD for continuous integration and deployment. This capability makes it a versatile option for teams that use GitHub for source control but prefer GitLab’s CI/CD features. Moreover, GitLab CI/CD supports extensive customization, making it ideal for projects that require intricate workflows and detailed stage definitions.
@@ 3. Study setup
This section presents our study setup. We explain how we collect and prepare the data for our studied RQs.
@@ 3.1. Data Collection
Figure 2 provides an overview of our data collection process. We focus on Android applications given the extensive research on them in the literature and the availability of public and open-source data. For data collection, we rely on two primary sources that have been commonly used by prior research: GitHub [^3] https://github.com and the F-Droid [^4] https://www.f-droid.org store (Cruz et al., 2019b; Lin et al., 2020; Pöll and Roland, 2022; Wang et al., 2023a).

### v2

@@ Abstract.
Continuous Integration and Continuous Delivery (CI/CD) is a well-established practice that automatically builds, tests, packages, and deploys software systems. To adopt CI/CD, software developers need to configure their projects using dedicated YML configuration files. Mobile apps have distinct characteristics with respect to CI/CD practices, such as testing on various emulators and deploying to app stores. However, little is known about the challenges and added value of adopting CI/CD in mobile apps and how developers maintain such a practice. In this paper, we conduct an empirical study on CI/CD practices in $2,557$ Android apps adopting four popular CI/CD services, namely GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD. We also compare our findings with those reported in prior research on general CI/CD practices to situate them within broader trends. We observe a lack of commonality and standardization across CI/CD services and Android apps, leading to complex YML configurations and associated maintenance efforts. We also observe that CI/CD configurations focus primarily on the build setup, with around half of the projects performing standard testing and only $9\%$ incorporating deployment. In addition, we find that CI/CD configurations are changed bi-monthly on average, with frequent maintenance correlating with active issue tracking, project size/age, and community engagement. Our qualitative analysis of commits uncovered $11$ themes in CI/CD maintenance activities, with over a third of the changes focusing on improving workflows and fixing build issues, whereas another third involves updating the build environment, tools, and dependencies. Our study emphasizes the necessity for automation and AI-powered tools to improve CI/CD processes for mobile apps and advocates creating adaptable open-source tools to efficiently manage resources, especially in testing and deployment.
@@
ccs: Software and its engineering Software configuration management and version control systemsccs: Software and its engineering Software deploymentccs: Software and its engineering Software development process managementccs: Information systems Mobile applications
@@ 1. Introduction
Continuous Integration and Continuous Delivery (CI/CD) is a well-established practice that automates the building, testing, packaging, and deployment of software systems (Fowler and Foemmel, 2006; Atlassian, [n.d.]; CircleCI, [n.d.]). CI/CD pipelines consist of automated processes designed to reduce human error, improve developer productivity, and accelerate release cycles. Their widespread adoption in both industry and open-source communities has led to extensive research on the challenges and benefits associated with CI/CD pipelines (Hilton et al., 2017, 2016a; Thatikonda, 2023).
Mobile applications, henceforth referred to as mobile apps, have become integral to our daily lives, serving various purposes, from communication to entertainment. Research reports that apps with more frequent release cycles tend to achieve higher ratings, as users often positively perceive regular updates and improvements (McIlroy et al., 2016). Research has also explored the benefits of adopting CI/CD in the software development lifecycle, notably enhancing quality and accelerating the delivery timeline. In the context of CI/CD, Cruz et al. (2019a) studied the testing practices in open-source mobile apps and found that the use of CI/CD pipelines is surprisingly limited among the mobile apps they analyzed. Mobile apps differ substantially from general-purpose software systems (e.g., desktop applications or web services) due to their unique ecosystem (Cruz et al., 2019a; Hu and Neamtiu, 2011; Picco et al., 2014). Developers must handle complex user interactions (e.g., swipe, pinch) (Zaeem et al., 2014) and account for devices with resource constraints, such as battery and lower processing power. Moreover, mobile apps need to support a wide range of operating system versions and hardware configurations (Khalid et al., 2014), which often causes traditional testing tools to fall short in this context (Wang and Alshboul, 2015; Maji et al., 2010). Given that mobile apps typically follow a tight weekly or bi-weekly release cycle (Nayebi et al., 2016), this can create pressure on testing and deployment tasks. In the context of CI/CD pipelines, while regular applications might run tests and deploy to cloud servers, mobile apps require additional steps that demand platform-specific tooling and plugins, such as multi-variant builds, app bundle generation (e.g., bundleRelease), emulator-based instrumentation testing (e.g., reactivecircus/android-emulator-runner), code signing (e.g., signingConfigs), and deployment to mobile platforms like the Google Play Store (e.g., actions/upload-artifact or fastlane). These tasks introduce distinct configuration challenges, which our study aims to investigate.
@@
Despite such importance, challenges in configuring and maintaining CI/CD for mobile apps have not been well studied. Though Liu et al. (2022) studied the adoption of CI/CD in mobile apps, their study mainly provided general statistical insights about CI/CD usage without exploring in-depth the specific configuration or maintenance practices of developers. In this paper, we conduct an empirical study to investigate the configuration practices and maintenance activities associated with CI/CD in a curated set of 2,557 open-source Android apps adopting four popular CI/CD services, namely GitHub Actions, Travis CI, CircleCI, and GitLab CI/CD. In particular, our study explores the common patterns and challenges that developers face with CI/CD configurations for their apps. To contextualize our observations within the broader trends of CI/CD, we compare and contrast our findings with previous research findings on general-purpose CI/CD practices. We aim to provide insights into the current state of CI/CD adoption within the Android development community and highlight areas for potential improvement. We address the following research questions (RQs).
RQ1: How do CI/CD configurations vary in complexity and tool usage for Android apps?
@@
In this RQ, we examine CI/CD configurations across multiple services. We observe that CI/CD configurations are not standardized, with complexities varying by service and type of configuration. We also find that most of the attention in CI/CD configuration is on the pipeline setup, accounting for more than 81% of the configurations. While testing is found to be present in about half of the projects, mainly through standard unit tests, deployment is comparatively less common, occurring in 9% of the projects. Our investigation shows that configuring deployment is notably complex and often does not directly target app stores, indicating a need for more integrated and automated solutions that streamline mobile app deployments to popular platforms, such as the Google Play Store.
@@
Adopting CI/CD in mobile app development makes developers responsible for the ongoing maintenance of their pipelines, which involves tasks such as feature additions and issue resolutions. In this RQ, we analyze and model the evolution of CI/CD configurations. We observe that developers regularly maintain CI/CD pipelines, yet not very frequently, mainly focusing on improving the setup and build phases. Over time, projects tend to simplify their CI/CD pipelines. While maintenance is found to occur bi-monthly on average, many projects update CI/CD pipelines more often, especially those using GitHub Actions. Our results show that increased maintenance activity correlates significantly with factors such as active issue tracking, project size/age, and community engagement.
@@
As CI/CD pipelines are continually updated, it is important to understand the rationale for these changes. In this RQ, we perform a manual analysis of a statistically significant random sample of commits related to CI/CD configurations. We observe that more than a third of CI/CD configuration changes focus on improving workflow and fixing build issues, while another third targets updates to the build environment, tools, and dependencies. To validate our manually identified themes for all commits in our dataset, we employ Latent Dirichlet Allocation (LDA) for automated topic generation and observe that the results are consistent with our observations.
@@
  • This study is the first of its kind in analyzing CI/CD configurations in Android apps across various services, using a dataset of 2,557 projects. It highlights current challenges in CI/CD adoption due to non-standardized practices.
  • This study models the evolution of CI/CD configurations using linear regression and identifies the potentially relevant factors related to changes in CI/CD pipelines.
  • This study presents a manually validated classification of 11 themes related to the maintenance of CI/CD pipelines, corroborated by automatically generated topics using LDA, ensuring both depth and breadth in our analysis.
  • This study contextualizes the CI/CD trends and practices observed in Android apps with prior findings on general-purpose CI/CD practices in open-source and industrial applications.
Overall, our study highlights the lack of standardized configurations in CI/CD pipelines for Android apps, emphasizing the need for uniform practices to improve interoperability. In addition, this study recommends employing automation tools and AI-driven solutions, such as bots and intelligent recommendation systems, to enhance and streamline CI/CD processes. Moreover, researchers are encouraged to develop adaptable frameworks and open-source tools to address resource constraints in mobile apps, especially those concerned with testing and deployment phases.
Paper organization: The rest of this paper is organized as follows. Section 2 provides background information relevant to our study. Section 3 details our study setup to investigate the challenges in configuring CI/CD pipelines for mobile apps. Section 4 outlines our research questions, providing both their motivation and the obtained results. We discuss the implications of our work in Section 5. Section 6 illustrates the potential validity threats of our findings. We present the related work in Section 7. Finally, we conclude our work and provide recommendations for future work in Section 8.
@@ 2. Background
Continuous Integration (CI) and Continuous Delivery (CD) are essential practices in modern software development, particularly for Android mobile apps (Tak and Modi, 2018). These practices ensure that code changes are automatically built, tested, and prepared for release, allowing for faster and more reliable app updates.
Continuous integration allows developers to frequently commit code changes to a shared, central repository, triggering automated build and test processes for smooth integration (Fowler and Foemmel, 2006). This helps catch bugs early, streamline new features, and improve code quality. Continuous delivery extends CI by automating deployment and preparing code changes for release (Humble and Farley, 2010). The deployment process in CI/CD involves moving integrated and tested changes to a production environment. Deployments can be manual or automated, depending on project needs. Automated deployments reduce the risk of human error and ensure consistency across different environments. Many CI/CD service providers (referred to as CI/CD services) are available to streamline the development process for Android apps. Examples include: GitHub Actions [^1] https://docs.github.com/en/actions , Travis CI [^2] https://docs.travis-ci.com, CircleCI [^3] https://circleci.com/docs, and GitLab CI/CD [^4] https://docs.gitlab.com/ci. Each service offers unique features and configurations tailored to various project needs and development workflows. Triggering CI/CD events can be done in several ways. Some services, like GitHub Actions, have native integrations that allow seamless triggering of CI/CD events within the same platform. Other services, like CircleCI, enable CI/CD builds to be triggered on third-party servers using webhooks for specific events, such as when code is pushed or a pull request is created. In this section, we describe the CI/CD process in general, how it is configured in the context of Android mobile app development, and the differences in CI/CD configuration across various services.
@@ 2.1. CI/CD Process
In an Android CI/CD pipeline, various processes are involved to ensure seamless integration and delivery. Figure 1 gives an overview of the CI/CD process in general. These processes are typically configured in YML files and include: Environment Setup, Dependency Management, Compiling/Building, Packaging, Testing, Deployment, and Feedback. In Listing 1, we provide a sample YML configuration for a CI/CD pipeline (in GitHub Actions) and give a brief description of each process. The red text represents the CI/CD directives (keys that define specific instructions) that guide the build, blue text represents executable commands, black text represents parameter values, and green text represents comments.
Figure 1. General overview of the CI/CD process.
  • Environment Preparation. In this stage, the CI/CD service initializes the environment needed to build and test an Android app, which involves choosing the machine operating system, specifying the appropriate version of the Java Development Kit (JDK), setting up Android SDK [^5] https://developer.android.com/studio components, and configuring any necessary environment variables (see lines 5-12).
  • Dependency Management. In this stage, the dependencies required to build and run an Android app are installed, which typically involves running tasks for the Gradle [^6] https://gradle.org build management tool, for example, to download and cache dependencies specified in build files (see lines 14-15).
  • Building. In this stage, an Android app is compiled by running respective build tasks to compile the source code and generate executables and resources (see lines 17-18).
  • Packaging. In this stage, an Android app is packaged into a certain distribution format, such as the Android Package (APK) [^7] https://developer.android.com/studio/profile/apk-profiler or the Android App Bundle (AAB) [^8] https://developer.android.com/guide/app-bundle, which comprises executables with any required configurations and third-party libraries (see lines 20-25).
@@
  • Deployment. In this stage, the successfully built and tested application is deployed to a production environment, which typically involves uploading the APK file to a distribution platform, such as the Google Play Store [^9] https://play.google.com (see lines 30-33).
@@ 2.2.1. GitHub Actions
A CI/CD service that is directly integrated within GitHub, allowing developers to define their CI/CD workflows using YML files placed in the .github/workflows directory. This service is highly flexible, enabling developers to customize every step of the CI/CD process using job and step directives. Its integration with GitHub repositories facilitates the triggering of workflows based on various events (e.g., pushes, pull requests, and more). In addition, GitHub Actions provides access to a wide range of pre-built actions available in the GitHub Marketplace, which can be easily incorporated into workflows.
@@ 2.2.2. Travis CI
An independent CI/CD service that simplifies the CI/CD setup with its use of a .travis.yml file located in the root of the repository. It employs predefined directives, such as install, script, before_script, and after_success, making it straightforward to set up and use. This allows developers to quickly establish a CI/CD pipeline without needing to define every process explicitly. Travis CI integrates seamlessly with GitHub repositories, making it a popular choice for many open-source projects due to its simplicity and ease of use.
@@ 2.2.3. CircleCI
An independent CI/CD service that uses a circle.yml or config.yml file found in the .circleci/ directory. It provides predefined jobs like checkout, run, and deploy, which help structure the CI/CD process while allowing for the inclusion of custom scripts for more complex setups. CircleCI is known for its strong support for Docker, enabling containerized builds that ensure consistency across different environments [^10] https://circleci.com/docs/about-circleci. This makes CircleCI suitable for projects requiring advanced build and deployment configurations, such as Android mobile apps.
@@ 2.2.4. GitLab CI/CD
A fully integrated within the GitLab version control system, but can also be used within GitHub. It uses a .gitlab-ci.yml file to define the CI/CD pipeline. Similar to Travis CI, it features predefined processes such as before_script, script, and after_script, while also allowing developers to create custom processes and jobs for greater flexibility. The integration with GitHub repositories is straightforward, allowing developers to connect their GitHub projects to GitLab CI/CD for continuous integration and deployment. This capability makes it a versatile option for teams that use GitHub for source control but prefer GitLab’s CI/CD features. Moreover, GitLab CI/CD supports extensive customization, making it ideal for projects that require intricate workflows and detailed stage definitions.
@@ 3. Study setup
This section presents our study setup. We explain how we collect and prepare the data for our studied RQs. Our data, scripts, and raw results can be found in our replication package (Ghaleb et al., 2024).
@@ 3.1. Data Collection
Figure 2 provides an overview of our data collection process. We focus on Android apps given the extensive research on them in the literature and the availability of public and open-source data. For data collection, we rely on two primary sources that have been commonly used in prior research: GitHub [^11] https://github.com and the F-Droid [^12] https://www.f-droid.org store (Cruz et al., 2019b; Lin et al., 2020; Pöll and Roland, 2022; Wang et al., 2023a).